5. You need to expand production, either in the U.S or in Mexico. Currently, your average American workers produce 100 units per hour and Mexican workers produce only 16 units per hour. But your American workers will cost you $20 per hour and your Mexican workers cost you $3 per hour.

a. Graph as much of this information as possible as isocost and isoquant information?

b. Where should you expand production?

c. Looking ahead into the near future, you forecast that your American workers will cost you $25 per hour, but demand for labor in mexico will drive up costs per worker there to $5 per hour. How does this alter the isocost and the isoquant graph?

d. Given these forecasts, where should you expand production?

Explain all answers.
Answer #1

A)let x = American worker

Y=Mexico worker

iso cost line equation,

C=20X+3Y

Take isoquant of Q =1600

B) as you can see in previous graph that iso cost line is tangent with isoquant line on Y axis ,which shows Mexico workers.,so firm should expand production in Mexico.

The per unit production from American worker=20/100=0.2

The per unit production cost form Mexico worker=3/16=0.18

So production cost is lower in Mexico ,so to minimize the cost and maximize the profit firm should expand production in mexico.

C)new iso cost lien equation,

C=25X+5Y

D) as you can see new iso cost line is tangent with same isoquant line at x axis which shows American worker,so firm should expand its production in America.

Per unit cost in America=25/100=0.25

Per unit cost in Mexico=5/16=0.31

So production cost is lower in America so firm should expand its production in America